# From Specificity to Generality: Revisiting Generalizable Artifacts in Detecting Face Deepfakes

### v1

@@ Abstract
Corresponding Author.
Detecting deepfakes has been an increasingly important topic, especially given the rapid development of AI generation techniques. In this paper, we ask: How can we build a universal detection framework that is effective for most facial deepfakes? One significant challenge is the wide variety of deepfake generators available, resulting in varying forgery artifacts (e.g., lighting inconsistency, color mismatch, etc). But should we “teach” the detector to learn all these artifacts separately? It is impossible and impractical to elaborate on them all. So the core idea is to pinpoint the more common and general artifacts across different deepfakes. Accordingly, we categorize deepfake artifacts into two distinct yet complementary types: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). FIA arise from the challenge of generating all intricate details, inevitably causing inconsistencies between the complex facial features and relatively uniform surrounding areas. USA, on the other hand, are the inevitable traces left by the generator’s decoder during the up-sampling process. This categorization stems from the observation that all existing deepfakes typically exhibit one or both of these artifacts. To achieve this, we propose a new data-level pseudo-fake creation framework that constructs fake samples with only the FIA and USA, without introducing extra less-general artifacts. Specifically, we employ a super-resolution to simulate the USA, while design a Blender module that uses image-level self-blending on diverse facial regions to create the FIA. We surprisingly found that, with this intuitive design, a standard image classifier trained only with our pseudo-fake data can non-trivially generalize well to unseen deepfakes.
@@ 1 Introduction
Face Xray [44], PCL [93], OST [4]
SLADD [3]
SBI [67]
SeeABLE [41]
RBI  [68]
@@
Table 1: Comparison of our framework and Sota methods using pseudo-deepfake synthesis. Our method differs from previous methods in terms of the level at which alterations are applied (local vs global), the introduced artifacts, and the level at which facial inconsistencies are applied.
In recent years, A growing number of facial manipulation techniques have advanced due to the rapid development of generative models [39, 40, 62, 80], which facilitate the production of highly realistic and virtually undetectable alterations. As a result, the risks of personal privacy being violated and the spread of misinformation have increased significantly. Therefore, developing a universal deepfake detector that can be used to detect the existing diverse fake methods has become an urgent priority.
In this paper, we pose the question: How can we build a universal detection framework that is effective for most facial deepfakes? A significant challenge lies in the wide range of deepfake generators, which lead to different types of forgery artifacts such as lighting inconsistency and color mismatch. Specifically, a large number of existing works are dedicated to creating detection algorithms that are carefully designed to identify counterfeit artifacts within specific handcrafted designs, including eye blinking [46], pupil morphology [26], and corneal specularity [33]. However, these methods are mainly effective for identifying particular forgery techniques. But should we train the detector to learn each of these artifacts separately? It is impossible and impractical to cover them all. So, where should we start from?
The core idea is to encourage the detection model to learn the more generalizable artifacts. As indicated in previous works [84, 67, 75], different fakes might share some common forgery patterns, and the unlimited deepfake artifacts could be generalized and summarized by the limited generalizable artifacts. Motivated by this, we conduct an in-depth preliminary exploration of the common patterns of existing deepfake artifacts. After our preliminary exploration, we identify two generalizable deepfake artifacts and categorize them into two distinct yet complementary categories: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA), as depicted in Figure 1. FIA represents an inherent limitation that deepfakes struggle to overcome, originating from the generator’s incapacity to precisely reproduce facial attributes and nuances, as well as the inescapable disparities between regions induced by amalgamation processes. In contrast to FIA, USA is not readily apparent to the unaided eye and arises from an up-sampling process that the generator’s decoder cannot adequately substitute. Considering the intrinsic necessity for decoding and up-sampling within the generator’s operations, every facial manipulation method invariably introduces Up-Sampling Artifacts (USA). Numerous prior investigations have substantiated this observation [22, 70]. To date, all existing deepfakes characteristically display one or both of these artifact categories.
To enable the model to simultaneously capture FIA and USA, we propose a sophisticated data augmentation method: FIA-USA, which leverages super-resolution models [76] and autoencoders [61] to introduce Up-Sampling Artifacts (USA) by reconstructing the face, and employs a strategy of generating multiple masks to blend the reconstructed face with the original, thereby introducing Face Inconsistency Artifacts (FIA) by creating discrepancies at both the facial and regional levels. Augmented with the aforementioned dual artifacts, our method effectively conditions the model to generalize across unseen deepfakes. Furthermore, to maximize the efficacy of FIA-USA, we introduce two complementary technique: Automatic Forgery-aware Feature Selection (AFFS), which streamlines feature dimensionality and augments feature discernment through adaptive selection. Region-aware Contrastive Regularization (RCR), by juxtaposing the features of the manipulated region against those of the authentic region, RCR enhances the model’s focus on the upsampling traces within the manipulated areas and the inconsistencies between different regions.
Our method surpasses the most recent state-of-the-art data augmentation methods [44, 67, 79, 86, 68, 43] on traditional deepfake benchmarks, and we have realized substantial improvements in generation deepfake benchmarks when compared to the baseline method. In addition, we perform cross-manipulation and cross-dataset detection evaluation on unseen data to align with real-world detection scenarios. Our method has shown enhanced generalization and robustness in the context of cross dataset generalization.
@@ 2 Related Work
Deepfake Generation Methods.
In the field of forgery detection, the broad definition of deepfake detection encompasses both forged face image / video detection and generated image / video detection. In this paper, deepfake detection specifically refers to the detection of fake faces. Based on the methods by which fake faces are generated, we categorize deepfakes into two types: traditional deepfake, and the generative deepfake. Traditional deepfake methods typically involve face replacement [47, 13, 18] and face reenactment [72, 73]. Facial replacement typically involves identity exchange using autoencoder-based [13] or graph based exchange methods [18], while facial reenactment employs reconstruction techniques to exchange faces from the source video to the target video while maintaining consistency in the identity of the target person. Unlike traditional deepfakes that involve local tampering, generative deepfakes directly generate fake faces through generative models [38, 59, 61, 37]. As discussed in some works [87, 86], the detection methods using RGB data augmentation methods have poor performance in this type of forgery, as these generation methods do not involve mixing operations. More importantly, an increasing number of facial tampering methods [40, 39, 40, 80] are evolving based on generation models, which implies that the fake faces generated by these new tampering methods will contain both the artifacts introduced by the generative models and the blend boundaries brought by the blend operations. To date, Most of the work only focuses on one forgery method and is ineffective for other forgery methods.
Classic Deepfake Detection.
Classic deepfake detection can be categorized into multiple aspects: On the one hand, deepfake image detection, which can also be applied to deepfake video detection by computing the average value of multiple frames. Typical methods encompass data augmentation [44, 67, 86, 93], frequency-domain analysis [60, 36, 53, 24], identity information [17, 9, 74], contrastive learning [23, 91], network structure improvement [92, 17], watermark [89, 77], robustness [42, 55], interpretability [16, 66] and so forth. On the other hand, deepfake video detection innovative video representation formats [83], video-level data augmentation [79], self-supervised learning [2, 29, 96], anomaly detection [19, 20], biological signal [81, 8], multi-modal based [58]. In addition to the above two aspects, the the remainder of the work involves forgery localization [55, 27], multi task learning [55, 90], real scenario [69],adversarial attack [48], and so on.
@@ 3.1 Face Inconsistency Artifacts (FIA)
As depicted in Figure 1, Face Inconsistency Artifacts (FIA) frequently occur in deepfakes due to the inherent difficulty in rendering intricate details, which often results in discrepancies between intricate facial features and the more uniform surrounding regions. In essence, until deepfake technology advances to the point where it can flawlessly replicate real faces, these artifacts are likely to persist. We will now explore the origins of Face Inconsistency Artifacts by categorizing deepfake techniques into two domains: Full Face Synthesis and Face Swapping. In the case of full face synthesis deepfakes, the presence of Face Inconsistency Artifacts is primarily contingent upon the capabilities of the generator, which, regardless of their magnitude, are invariably present. For face swapping, these artifacts arise from two primary sources: 1). the efficacy of the generator. 2). the blending seam created during the face swapping’s blending procedure. Furthermore, we can categorize face swapping into two main types based on the forged area: macro-editing and micro-editing. Macro-editing can be further classified into two subtypes: editing based on 4 facial keypoints and editing based on 81 keypoints, whereas micro-editing allows for precise adjustments to each facial feature. In section B of the appendix, we provide a detailed description of the classification of forgery techniques and their differences.
3.2 Up-Sampling Artifacts (USA)
@@
Conversely, the artifacts invariably present in the forged regions produced by the forgery generator are due to the upsampling process within the generator’s decoder, known as Up-Sampling Artifacts (USA). The extent of USA introduced can differ among various generators. To investigate more universal upsampling artifacts, we categorize forgery techniques into two types: One category generates images of forged regions that are as clear and free of blur as the original image (using Diffusion models [61] and GANs [38] as the main methods, and upsampling the latent code through autoencoder (AE) to generate images). And another type result in blurred tampered areas, requiring additional facial super-resolution models (SR) to process the blurred areas.
3.3 FIA-USA
We propose a new data augmentation method to generate pseudo-fake training samples, aiming to introduce FIA and USA present in forged faces. The core idea is to introduce FIA and USA simultaneously by blending the real image $I_{R}$ and the reconstructed image $I_{F}$, as follows:
$I_{PF}=M\odot I_{F}+(1-M)\odot I_{R},$ (1)
where $\odot$ represents the element wise multiplication operator, $M$ represents the mask of the manipulated area, which varies between 0 and 1. We considered two ways of generating masks to cover all possible occurrences of FIA: 1) Mask erosion on the convex hull of the facial 81 or 4 keypoints, aimed at simulating macro-editing; 2) Random combination of facial features, aimed at simulating micro-editing. Consequently, we address generic USA through three approaches: 1. Employ an autoencoder (AE) to reconstruct the entire image [61]. 2. Enhance the facial details using a super-resolution model [76] (SR). 3. Do not reconstruct the original image. Formally speaking, we introduced different levels of FIA through blending operations and reconstructed images, while introducing USA for forged areas through reconstructed images. Please refer to  Algorithm 1 for the complete algorithmic process. Table 1 compares the differences between our proposed FIA-USA method and the previous blend-based data augmentation method. In section C of the appendix, we provide more information on blend-based data augmentation and FIA-USA.
@@ 3.3 FIA-USA
1:def $\mathcal{T}$($I$) : $\triangleright$ Source-Target Augmentation
@@
4:     return $I$
5:def $Recon$($I$) : $\triangleright$ Reconstruct the source image
@@
11:         $I_{s}\leftarrow I$
12:     return $I$
13:$I_{s},I_{t}\leftarrow Recon(I),I$
@@
18:$M\leftarrow\text{CombineFacialFeatures}(L)$ or $\text{ConvexHull}(L)$
19:$M\leftarrow\text{MaskAugmentation}(M)$
20:$I_{PF}\leftarrow I_{s}\odot M+I_{t}\odot(1-M)$
Figure 2: The overall framework of our proposed method consists of: 1) FIA-USA, which categorizes deepfake artifacts into Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). By employing data augmentation techniques, the model is enhanced to concurrently recognize and learn from both categories of artifacts. 2) Region-aware Contrastive Regularization (RCR), which enables the model to focus simultaneously on forged boundaries and spatial artifacts through the contrastive regularization of features from forged and real regions. 3) Automatic Forgery-aware Feature Selection (AFFS), which achieves efficient feature selection by assessing the similarity between the feature maps of each channel in each layer and the mask.

### v2

@@ Abstract
Detecting deepfakes has been an increasingly important topic, especially given the rapid development of AI generation techniques. In this paper, we ask: How can we build a universal detection framework that is effective for most facial deepfakes? One significant challenge is the wide diversity of existing deepfake generators, which produced varied types of forgery artifacts (e.g., lighting inconsistency, color mismatch, etc). But should we “teach” the detector to learn all these artifacts separately? It is impossible and impractical to elaborate on them all. So the core idea is to pinpoint the more common and general artifacts across different deepfakes. Through systematic analysis of shared technical frameworks in existing deepfake algorithms, we categorize deepfake artifacts into two distinct yet complementary types: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). FIA arise from the challenge of generating all intricate details, inevitably causing inconsistencies between the complex facial features and relatively uniform surrounding areas. USA, on the other hand, are the inevitable traces left by the generator’s decoder during the up-sampling process, with all existing deepfakes exhibiting either or both artifacts. Subsequently, we propose a novel image-level pseudo-fake creation framework that constructs fake samples with only the FIA and USA, without introducing extra less-general artifacts. Specifically, we reconstruct the target face to simulate the USA, while utilize image-level blend on diverse facial regions to create the FIA. We surprisingly found that, with this intuitive design, a standard image classifier trained only with our pseudo-fake data can non-trivially generalize well to unseen deepfakes.
@@ 1 Introduction
In recent years, A growing number of facial manipulation techniques have advanced due to the rapid development of generative models kawar2023imagic ; kim2022diffface ; rosberg2023facedancer ; wu2023latent , which facilitate the production of highly realistic and virtually undetectable alterations. As a result, the risks of personal privacy being violated and the spread of misinformation have increased significantly. Therefore, developing a universal deepfake detector that can be used to detect the existing diverse fake methods has become an urgent priority.
In this paper, we pose the question: How can we build a universal detection framework that is effective for most facial deepfakes? A significant challenge lies in the wide range of deepfake generators, which lead to different types of forgery artifacts. Specifically, a large number of existing works are dedicated to creating detection algorithms that are carefully designed to identify counterfeit artifacts within specific handcrafted designs, including eye blinking li2018ictu , pupil morphology guo2022eyes , and corneal specularity hu2021exposing . However, these methods are mainly effective for identifying particular forgery techniques. But should we train the detector to learn each of these artifacts separately? It is impossible and impractical to cover them all. So, where should we start from?
@@
The core idea is to encourage the detection model to learn the more generalizable artifacts. As indicated in previous works yan2023ucf ; shiohara2022detecting ; wang2020cnn , different fakes might share some common forgery patterns, and the unlimited deepfake artifacts could be generalized and summarized by the limited generalizable artifacts. Motivated by this, we conduct an in-depth preliminary exploration of the common patterns of existing deepfake artifacts, through which we identify two generalizable deepfake artifacts and categorize them into two distinct yet complementary categories: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA), as depicted in Figure 1. FIA represents an inherent limitation that deepfakes struggle to overcome, originating from the generator’s incapacity to precisely reproduce facial attributes and nuances, as well as the inescapable disparities between regions induced by amalgamation processes. Unlike FIA, Up-Sampling Artifacts (USA) are not readily apparent to the unaided eye and originate from the generator’s inherent up-sampling process that the generator’s decoder cannot adequately substitute. Crucially, since decoding and up-sampling constitute fundamental operations in facial manipulation generators, such artifacts become an inevitable byproduct across all methods in this domain. Numerous prior investigations have substantiated this observation frank2020leveraging ; tan2024rethinking . To date, all existing deepfakes characteristically display one or both of these artifact categories.
To enable the model to simultaneously capture FIA and USA, we propose a sophisticated data augmentation method: FIA-USA, which leverages super-resolution models wang2021gfpgan and autoencoders rombach2022high to introduce Up-Sampling Artifacts (USA) by reconstructing the face, and employs a strategy of generating multiple masks to blend the reconstructed face with the original, thereby introducing Face Inconsistency Artifacts (FIA) by creating discrepancies at both the facial and regional levels. Augmented with the aforementioned dual artifacts, our method effectively conditions the model to generalize across unseen deepfakes. Furthermore, to maximize the efficacy of FIA-USA, we introduce two complementary technique: Automatic Forgery-aware Feature Selection (AFFS), which streamlines feature dimensionality and augments feature discernment through adaptive selection. Region-aware Contrastive Regularization (RCR), by juxtaposing the features of the manipulated region against those of the authentic region, RCR enhances the model’s focus on the upsampling traces within the manipulated areas and the inconsistencies between different regions.
Extensive experiments conducted on seven widely used Deepfake detection datasets (encompassing over 58 distinct forgery methods, spanning facial manipulation categories including identity swapping, expression reenactment, generative face synthesis, and attribute editing) validate the superior efficacy of our method. Detailed ablation experiments and robustness experiments have demonstrated the effectiveness of each component in our method and the robustness of our method to disturbances.
Table 1: Comparison of our framework and state-of-the-art(SOTA) methods using pseudo-deepfake synthesis. Our method differs from previous methods in terms of the level at which alterations are applied (local vs global), the introduced artifacts, and the level at which facial inconsistencies are applied.
@@
Face Xray li2020face , PCL zhao2021learning , OST chen2022ost
SLADD chen2022self
SBI shiohara2022detecting
SeeABLE larue2023seeable
RBI  sun2023generalized
Plug-and-Play yan2024generalizing
@@ 2 Related Work
Classic Deepfake Detection. Classic deepfake detection can be categorized into multiple aspects: On the one hand, typical deepfake image detection methods encompass frequency-domain analysis qian2020thinking ; luo2021generalizing , identity information dong2022protecting ; cozzolino2021id ; tinsley2021face , contrastive learning fung2021deepfakeucl ; zhang2023contrastive , network structure improvement zhao2021multi ; dong2022protecting , watermark yu2020responsible ; wang2022deepfake , robustness le2023quality ; nguyen2024laa , interpretability dong2022explaining ; shao2022detecting and so forth. On the other hand, deepfake video detection methods include innovative video representation formats xu2023tall , self-supervised learning cao2022end ; haliassos2022leveraging ; zhuang2022uia , anomaly detection fei2022learning ; feng2023self , biological signal wu2024local ; ciftci2020fakecatcher , multi-modal based oorloff2024avff . In addition to the above two aspects, the the remainder of the work involves forgery localization nguyen2024laa ; guo2023hierarchical , multi task learning nguyen2024laa ; yuan2022forgery , adversarial attack liang2024poisoned , and so on.
Deepfake Detectors Based on Data Synthesis. A notable approach involves synthesizing pseudo-fake faces during training to enhance generalization capability. Early works like Face X-ray li2020face pioneered blending-based augmentation by fusing facial regions from different identities to simulate forgery boundaries. Subsequent methods improved upon this paradigm through different artifact simulation strategies: SLADD chen2022self focused on local adversarial perturbations, SBI shiohara2022detecting proposed self-blending to amplify facial inconsistencies, SeeABLE larue2023seeable which proposes fine-grained region-specific blending, and Plug-and-Play yan2024generalizing employs facial feature masks rather than full-face manipulation for enhanced face synthesis precision. However, as highlighted in Table 1, existing methods predominantly focus on isolated artifact types (Mainly focused on FIA) and single-scale modifications (global or local), fundamentally limiting their ability to capture the artifacts interplay inherent in real deepfakes. The paradigm of this data augmentation method can be represented as:
$I_{F}=M\odot I_{t}+(1-M)\odot I_{s},$ (1)
where, $I_{t}$ represents the target face, $I_{s}$ represents the source face, and $M$ represents the mask. Our proposed FIA-USA method injects new vitality into this paradigm by generating multiple types of masks, introducing USA to $I_{t}$, as well as random combination of artifacts.
@@ 3 Method
Figure 2: The overall framework of our proposed method consists of: 1) FIA-USA enhances the capability of standard classifiers to effectively detect and generalize to unknown deepfake techniques by generating pseudo-fake data containing Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). 2) Region-aware Contrastive Regularization (RCR) enables the model to focus simultaneously on forged boundaries and spatial artifacts through the contrastive regularization of features from forged and real regions. 3) Automatic Forgery-aware Feature Selection (AFFS) achieves efficient feature selection by assessing the similarity between the feature maps of each channel in each layer and the mask.
@@
Face Inconsistency Artifacts (FIA). As depicted in Figure 1, Face Inconsistency Artifacts (FIA) frequently occur in deepfakes due to the inherent difficulty in rendering intricate details, which often results in discrepancies between intricate facial features and the more uniform surrounding regions. In essence, until deepfake technology advances to the point where it can flawlessly replicate real faces, these artifacts are likely to persist. We explore the origins of Face Inconsistency Artifacts by categorizing deepfake techniques into two domains: Full Face Synthesis and Face Swapping. In deepfake-based full face synthesis, Face Inconsistency Artifacts inherently persist regardless of the generator’s capabilities, while their severity directly correlates with the model’s performance. For face swapping, these artifacts arise from two primary sources: 1). the efficacy of the generator. 2). the blending boundary created during the face swapping’s blending procedure. Furthermore, we can categorize face swapping into two main types based on the forged area: a. macro-editing realize face swapping through 4 or 81 facial keypoints. and b. micro-editing, which enables granular control over individual facial features. In section B of the appendix, we provide a detailed description of the classification of forgery techniques and their differences.
@@
Up-Sampling Artifacts (USA). These artifacts consistently present in synthetic facial regions originate from upsampling operations within the generator’s decoder architecture, where interpolation processes inevitably introduce characteristic distortion patterns, with the extent of USA (Up-Sampling Artifacts) introduced varying significantly across different generator implementations. To investigate more universal up-sampling artifacts, we categorize forgery techniques into two types: One category generates images of forged regions that are as clear and free of blur as the original image. And another type result in blurred tampered areas, requiring additional facial super-resolution models (SR) to process the blurred areas.
FIA-USA. While FIA and USA represent two fundamental artifact categories in deepfakes, conventional data augmentation methods predominantly focus on simulating isolated forgery traces, thereby hindering detection models from effectively learning cross-domain forgery patterns. To bridge this critical gap, we propose FIA-USA—a dual-artifact collaborative enhancement framework engineered for universal detection. Our framework systematically emulates the forgery process through a three-pronged methodology:
1. Multi-Type Mask Synthesis(MTMS): The Multi-Type Mask Synthesis module aims to comprehensively model Face Inconsistency Artifacts (FIA) through a hierarchical mask generation strategy, covering both macro-editing boundaries and micro-editing traces. This process involves two complementary approaches:
a. Macro-Editing Mask. We generate facial contours through two configurations: (1) a high-precision mode using 81 facial keypoints to compute detailed convex hull boundaries, and (2) a rectangular approximation mode derived from 4 keypoints. The initial mask $\mathbb{M}$ is formulated as:
$\mathbb{M}=\mathcal{H}(K_{n}),n\in\{4,81\}$ (2)
where $\mathcal{H}(\cdot)$ denotes convex hull computation and $K_{n}$ represents the keypoints set. As same as shiohara2022detecting , the initial mask $\mathbb{M}$ subsequently undergoes erosion or dilation by applying two Gaussian filters with distinct kernel sizes, where erosion occurs when the kernel of the first filter is larger than that of the second, while dilation manifests when the first kernel is comparatively smaller.
b. Micro-Editing Masks. Given a set of 81 facial keypoints $K_{81}=\{k_{1},k_{2},\dots,k_{81}\}$, we first extract two subsets: eyes-specific keypoints $K_{eyes}\subset K$ and mouth-specific keypoints $K_{mouth}\subset K.$ Using convex hull computation $\mathcal{H}(\cdot)$, we derive two binary masks: the eye mask $M_{e}=\mathcal{H}(K_{eyes})$ and the mouth mask $M_{m}=\mathcal{H}(K_{mouth})$. These masks are then logically combined through union operations $\cup$ to generate three distinct facial regions: (1) $M=M_{e}$, (2) $M=M_{m}$, and (3) $M=M_{e}\cup M_{m}$ (combined mouth and eyes). Since other facial regions such as eyebrows and nose are rarely targeted in localized facial manipulations (e.g., attribute editing, expression synthesis), these features were intentionally excluded from our mask generation framework to prioritize regions most susceptible to forgery (eyes and mouth).
2. Multi-Modal Reconstruction(MMR): We employ two complementary reconstruction paradigms to reconstruct the target face($I_{t}$) in Formula  1, thereby systematically simulating the Up-Sampling Artifacts (USA) inherent in diverse deepfake generators: a. Autoencoder (AE) Reconstruction: Reconstruct $I_{t}$ through AE rombach2022high to simulate the excessive dependency relationship between adjacent pixels tan2024rethinking caused by upsampling in GAN / Diffusion models. b. Super-Resolution (SR) Reconstruction: Applies SR models wang2021gfpgan to upsample $I_{t}$, thereby replicating the characteristic artifacts inherent in deepfake post-processing pipelines. Specifically, given an input face image $I$, we generate reconstructed versions $I_{AE}=AE(I)$ and $I_{SR}=SR(I)$.
[…]
As shown in Table 1, compared to existing methods, FIA-USA has achieved collaborative artifact enhancement across local / global, facial / regional levels. Please refer to Algorithm 1 for the complete algorithmic process.
@@
1:def $\mathcal{T}$(II) : $\triangleright$ Source-Target Augmentation
@@
4:  return II
5:def ReconRecon(II) : $\triangleright$ Reconstruct the source image
@@
11:   $I_{t}\leftarrow I$
12:  return II
[…]
@@
[…]